UNIVERSIT. . AT BONN Physikalisches Institut

Total Cross Section Measurements in γp and γ^{∗}p at HERA

by Adolf Bornheim

Scale Issues in High-Energy Diffractive Vector-Meson Production

by James A. Crittenden

Contributions to the proceedings of the LISHEP International School on High-Energy Physics, Rio de Janeiro, Brazil, 16-20 February 1998

Post address: Nussallee 12 53115 Bonn Germany BONN-HE-98-02 Bonn University June 1998

ISSN-0172-8733